# Direct Preference Optimization: Your Language Model is Secretly a Reward Model

### v1

@@ A.1 Deriving the Optimum of the KL-Constrained Reward Maximization Objective
$\min_{\pi}\mathbb{E}_{x\sim\mathcal{D}}\left[\mathbb{D}_{\text{KL}}(\pi(y|x)\mid\mid\pi^{*}(y|x))+Z(x)\right]$ (14)
@@ Appendix B DPO Implementation Details and Hyperparameters
Unless noted otherwise, we use a $\beta=0.1$, batch size of 64 and the Adam optimizer with a learning rate of 1e-6 by default. We linearly warmup the learning rate from 0 to 1e-6 over 150 steps. For TL;DR summarization, we use $\beta=0.5$, while rest of the parameters remain the same.
@@ Appendix C Further Details on the Experimental Set-Up
C.1 GPT-4 prompts for computing summarization and dialogue win rates
A key component of our experimental setup is GPT-4 win rate judgments. In this section, we include the prompts used to generate win rates for the summarization and dialogue experiments. The order of summaries or responses are randomly chosen for every evaluation.
@@ C.1 GPT-4 prompts for computing summarization and dialogue win rates
C.2 Unlikelihood baseline
While we include the unlikelihood baseline (simply maximizing $\log p(y_{w}|x)$, the log probability of the preferred response, while minimizing $\log p(y_{l}|x)$, the log probability of the dispreferred response) in our sentiment experiments, we do not include it as a baseline in either the summarization or dialogue experiment because it produces generally meaningless responses, which we believe is a result of unconstrained likelihood minimization.
@@ D.2 Sample Responses and GPT-4 Judgments
In this section, we present examples of comparisons between DPO and the baseline (PPO temp 0. for summarization, and the ground truth chosen response for dialogue). See Tables 3-5 for summarization examples, and Tables 6-9 for dialogue examples.

### v2

@@ A.1 Deriving the Optimum of the KL-Constrained Reward Maximization Objective
$\min_{\pi}\mathbb{E}_{x\sim\mathcal{D}}\left[\mathbb{D}_{\text{KL}}(\pi(y|x)\mid\mid\pi^{*}(y|x))-\log Z(x)\right]$ (14)
@@ Appendix B DPO Implementation Details and Hyperparameters
Unless noted otherwise, we use a $\beta=0.1$, batch size of 64 and the RMSprop optimizer with a learning rate of 1e-6 by default. We linearly warmup the learning rate from 0 to 1e-6 over 150 steps. For TL;DR summarization, we use $\beta=0.5$, while rest of the parameters remain the same.
@@ Appendix C Further Details on the Experimental Set-Up
C.1 IMDb Sentiment Experiment and Baseline Details
The prompts are prefixes from the IMDB dataset of length 2-8 tokens. We use the pre-trained sentiment classifier siebert/sentiment-roberta-large-english as a ground-truth reward model and gpt2-large as a base model. We use these larger models as we found the default ones to generate low-quality text and rewards to be somewhat inaccurate. We first use supervised fine-tuning on a subset of the IMDB data for 1 epoch. We then use this model to sample 4 completions for 25000 prefixes and create 6 preference pairs for each prefix using the ground-truth reward model. The RLHF reward model is initialized from the gpt2-large model and trained for 3 epochs on the preference datasets, and we take the checkpoint with the highest validation set accuracy. The “TRL” run uses the hyper-parameters in the TRL library. Our implementation uses larger batch samples of 1024 per PPO step.
C.2 GPT-4 prompts for computing summarization and dialogue win rates
A key component of our experimental setup is GPT-4 win rate judgments. In this section, we include the prompts used to generate win rates for the summarization and dialogue experiments. We use gpt-4-0314 for all our experiments. The order of summaries or responses are randomly chosen for every evaluation.
@@ C.2 GPT-4 prompts for computing summarization and dialogue win rates
C.3 Unlikelihood baseline
While we include the unlikelihood baseline [44] (simply maximizing $\log p(y_{w}|x)$, the log probability of the preferred response, while minimizing $\log p(y_{l}|x)$, the log probability of the dispreferred response) in our sentiment experiments, we do not include it as a baseline in either the summarization or dialogue experiment because it produces generally meaningless responses, which we believe is a result of unconstrained likelihood minimization.
@@ D.2 Sample Responses and GPT-4 Judgments
In this section, we present examples of comparisons between DPO and the baseline (PPO temp 0. for summarization, and the ground truth chosen response for dialogue). See Tables 4-6 for summarization examples, and Tables 7-10 for dialogue examples.